# Thermal Comfort in Sight: Thermal Affordance and its Visual Assessment for Sustainable Streetscape Design

### v1

@@ 5.3 VATA mapping and spatial analysis
Figures 11 and 12 illustrate the spatial distribution of high and low VATA values, respectively. High-VATA areas, particularly city parks such as Garden of Peace, Windsor Nature Park, and Bukit Timah Nature Reserve, offer superior thermal comfort due to abundant green spaces, highlighting their importance for urban planning and public health. The exceptional VATA performance of the area around East Coast Park further underscores the cooling effects of water bodies and greenery in urban heat mitigation. In contrast, Figure 12 reveals low-VATA regions in the northwestern and eastern parts of Singapore, presenting opportunities for improving streetscape thermal affordance. Thus, the VATA framework serves as a valuable tool for urban planners and designers to identify and improve thermal affordance levels across different parts of a city, aiding the development of sustainable and livable environments.
5.4 IF-VPI-VATA two-stage inference modelling and feature weight analytics
@@ 5.4.1 Overview of two-stage ENRM model performance
Table 4 presents the results for IF-VATA, VPI-VATA, and two-stage IF-VPI-VATA model. The two-stage model achieved the best performance with an adjusted $R^{2}$ of 0.7443, comparable to the MTNNL-based prediction model (adjusted $R^{2}$ = 0.7316).
Figure 13: Sankey diagram of positive feature weight proportion in the IF-VPI-VATA two-stage inference modeling.
Figure 14: Sankey diagram of negative feature weight proportion in the IF-VPI-VATA two-stage inference modeling
5.4.2 Feature weight analytics in the two-stage modelling for thermal affordance design implications
@@ 6.1 VATA framework on sustainable streetscape planning and design
Figure 15: Long-term sustainable streetscape planning and design based on continous thermal affordance evaluation supported by VATA framework.
@@ 6.2 Limitations and future work
The study has limitations and opportunities for future research. Temporal variance in SVI was not considered, omitting the possible impacts of time, weather conditions, and seasons on thermal affordance assessment. Future research should sample street views from the same locations at various times to capture these influences. While visual cues provide valuable insights, they are insufficient for a comprehensive thermal affordance evaluation. Field research is crucial to bridge the gap between visual-based predictions and empirical evaluations on-site. Although in-field comfort data validates our VATA predictions, it is not used for the prediction or inference process. Future work should integrate visual assessments and field data to improve prediction accuracy and inference comprehension.
Another limitation of this study is that the validation was conducted along a single continuous path, which limits the diversity of environmental conditions captured. Although the 43 investigation points provide valuable data, they are confined to one specific route, which may not represent the full range of urban settings. To improve the robustness and generalisability of the model, future research should explore validation across more diverse field locations, covering a broader variety of urban environments.
The subjective visual assessments from online surveys may introduce bias due to participants’ diverse backgrounds, lacking professional expertise, or full representation of all city residents. Despite this limitation, online surveys offer a practical means of collecting data within manageable constraints. Future research should address these limitations by incorporating temporal variations, integrating field data for better predictions, and using more representative sampling methods for subjective evaluations.
@@ 7 Conclusions and Contributions
    Introducing thermal affordance, which describes the built environment’s hidden information and inherent capability to promote thermal comfort. This new concept integrates objective descriptions of fixed environmental impacts on thermal comfort perception.
@@
    Pioneering the VATA framework, which integrates objective IF from SVI with subjective VPI from survey data to build a two-stage neural network model for SVI-based thermal affordance assessment, offering a novel approach to OTC evaluation in urban settings. VATA mapping helps urban planners to prioritise regions for future streetscape sustainable renovation.
@@
    Exploring the IF-VPI-VATA relationship through inference modelling, which contributes to understanding the inference relationship between objective streetscape compositions, subjective human perceptions, and VATA, informing future sustainable streetscape planning by providing key logic chains on how streetscape compositions increase or decrease streetscapes’ thermal affordance.
@@
    Generating a geospatially precise urban-scale mapping of thermal affordance across Singapore’s streetscapes. This application of the VATA framework offers valuable insights for long-term evaluation, design, and implementation of sustainable urban streetscape solutions.
@@
    The validity of the VATA framework has been demonstrated by in-field comfort data, with comfort modelling based on VATA metrics outperforming models based solely on IF metrics. The inclusion of VPI in the two-stage modelling is the key factor behind this performance improvement.
@@ Appendix C Subfeatures of IF datasets in VATA modelling
Appendix D Explanation of the rationale behind the training sample size
Figures 17 demonstrate that a sample size of 300 images is sufficient to estimate the average VATA score within a 95% confidence interval (CI), with an error margin of only 5%. This estimation assumes an ideal population mean of 2.5 (based on the normal distribution assumption of the TrueSkill algorithm) and a data-augmented population standard deviation of 1. Additionally, increasing the sample size does not result in a significant reduction in CI width, as shown by the graph illustrating the second derivative of the CI width. Moreover, as the sample size increases from 300 to 500, the distribution of VATA scores remains largely unchanged (K-S statistic $<$ 0.05), further suggesting that a representative and convergent sample of 300 images is convergent and sufficient for training a VATA prediction model.
Figure 17: Training set of 300 images is convergent on scoring distribution and sufficient for prediction model training.
@@ Appendix E Details on MTNNL model training data and process
Figures 18.a and 18.c demonstrate that different streetscape categories can influence their VATA scores obtained from the online survey. During the division of train/validation/test sets, it is ensured that an equal quantity of each of the five streetscape categories is included in every fold. Figure 18.b illustrates the training process for fold 1 in k-fold cross-validation model training. Both validation loss and train loss consistently decrease and stabilise, indicating that no overfitting occurred during training.
@@ Appendix F Details on stage-1 ENRM models
From the stage-1 ENRM models shown in Figure 6, we can observe that the $R^{2}$ values for IF-VPI inference modeling in the microclimate inference and environmental assessment sections—except for ‘material comfort’—exceed 0.6. This suggests that these VPIs are highly predictable and can be effectively modeled using IFs. Additionally, the VPIs ‘enclosure’, ‘transparency’, and ‘complexity’ in the design quality section demonstrate strong consistency, while in the evoked emotion section, only the $R^{2}$ for ‘lively’ exceeds 0.6. This indicates that most VPIs related to evoked emotions vary significantly between individuals. It is crucial to consider both IF and VPI dimensions in VATA evaluation, as visual-perceptual indicators capture human visual-perceptual interpretations that are not fully reflected in SVI image features.
@@
Appendix G Further analytics on featurewise relationships
G.1 Correlation among VATA and visual-perceptual indicators (VPI)
@@
G.2 Relationship between image features (IF) and visual-perceptual indicators (VPI)

### v4

@@ 4.3 VATA mapping and spatial analysis
Figures 11 and 12 illustrate the spatial distribution of high and low VATA values, respectively. High-VATA areas, particularly city parks such as Garden of Peace, Windsor Nature Park, and Bukit Timah Nature Reserve, offer superior thermal comfort due to abundant green spaces, highlighting their importance for urban planning and public health. The exceptional VATA performance of the area around East Coast Park further underscores the cooling effects of water bodies and greenery in urban heat mitigation. In contrast, Figure 12 reveals low-VATA regions in the northwestern and eastern parts of Singapore, presenting opportunities for improving streetscape thermal affordance. Thus, the VATA framework serves as a valuable tool for urban planners and designers to identify and improve thermal affordance levels across different parts of a city, aiding the development of sustainable and liveable environments.
4.4 IF-VPI-VATA two-stage inference modelling results
@@
Table 4 presents the results for IF-VATA, VPI-VATA, and two-stage IF-VPI-VATA model. The two-stage model achieved the best performance with an adjusted $R^{2}$ of 0.7443, comparable to the MTNNL-based prediction model (adjusted $R^{2}$ = 0.7316). The feature weights derived from the two-stage IF–VPI–VATA model highlight the positive and negative factors influencing thermal affordance. A detailed discussion of these factors, including positive features (e.g., vegetation, shading area, human scale) and negative features (e.g., traffic elements, complexity, sunlight intensity), as well as design insights drawn from the inference modelling, is provided in G.
@@ 5.1 VATA framework on sustainable streetscape planning and design
Importantly, while our current research focuses on the hot tropical climate of Singapore, the VATA framework and its underlying workflow can be methodologically transferred to other cities. To achieve reliable outcomes under different climatic and urban conditions, it is essential to conduct new local SVI visual assessment surveys and incorporate locally relevant geospatial data. Moreover, performing on-site thermal comfort validations in the target cities would be a crucial step to fine-tune the model’s parameters, ensuring the VATA approach remains robust and accurate. By doing so, the VATA workflow can guide sustainable streetscape planning and design across diverse environments while maintaining methodological consistency, adaptability, and scientific rigour.
Figure 13: Long-term sustainable streetscape planning and design based on continuous thermal affordance evaluation supported by VATA framework.
@@ 5.2 Limitations and future work
The primary objective of this research is to leverage visual features in SVI and human assessment for a cost-effective, scalable approach to thermal affordance evaluation. However, certain limitations suggest opportunities for future study. First, temporal variance in SVI was not considered, omitting the possible impacts of time, weather conditions, and seasons on thermal affordance assessment. Future research could address this by sampling street views from the same locations at various times. Second, the validation was conducted along a single continuous path with 43 investigation points, which limits the diversity of environmental conditions captured. Expanding to a broader selection of field locations and incorporating microclimatic parameters (e.g. humidity, air temperature, and wind speed) for the validation would enhance both model robustness and generalizability.
Additionally, while visual cues provide valuable insights, they represent only one facet of urban conditions. Future research might incorporate complementary data sources such as satellite imagery, remote sensing methods, or GIS-based analyses (e.g. land surface temperature or vegetation indices) to achieve a more comprehensive understanding of thermal affordance. Online surveys also pose limitations due to potential biases from participants’ varied backgrounds and limited professional expertise; however, they remain a practical means of data collection within manageable constraints.
Finally, testing and validating this approach in multiple urban contexts, possibly spanning diverse climatic zones and geospatial conditions, would refine the framework and bolster its applicability. Through these efforts, future work can offer a more robust, scalable, and universally relevant tool for assessing and improving thermal comfort in urban environments worldwide.
@@ 6 Conclusions and Contributions
    Introducing thermal affordance, which describes the built environment’s hidden information and inherent capability to promote thermal comfort. This new concept represents objective and fixed environmental influences on thermal comfort perception.
@@
    Pioneering the VATA framework, which integrates objective image features (IF) derived from street view imagery (SVI) with subjective visual-perceptual indicators (VPI) from survey data to build a two-stage neural network model for SVI-based cost-effective urban-scale thermal affordance assessment.
@@
    Exploring the IF–VPI–VATA relationship through inference modelling, contributing to the understanding of how visual compositions and human perceptions jointly shape thermal affordance, informing future sustainable streetscape planning and design to improve thermal comfort perception.
@@
    Generating a spatially detailed urban-scale mapping of thermal affordance across Singapore’s streetscape, which is validated by in-field thermal comfort data and transferable to other cities, prioritising regions for future sustainable streetscape renovation.
@@
    Demonstrating that the validity of the VATA framework is supported by in-field comfort data. Comfort modelling based on VATA metrics outperforms models relying solely on IF metrics, highlighting the key role of VPI in the two-stage modelling process.
@@
Data and Code Availability
The computed VATA data for thermal affordance in Singapore can be accessed at https://thermal-affordance.ual.sg/. Data for additional cities will be updated regularly. The code and calculation methods will be open-sourced at https://github.com/Sijie-Yang/Thermal-Affordance.
@@ Appendix C Subfeatures of IF datasets in VATA modelling
Appendix D Rationale explanation of the survey and training sample sizes
Given the constraints on resources and participant availability for the visual evaluation, it was essential to carefully determine an appropriate sample size. An excessively large number of images would risk incomplete evaluations, leading to inaccurate results, while too small a sample would fail to comprehensively represent the diversity of street scenes. Thus, a balanced approach was needed to ensure both feasibility and representativeness.
Our solution was to employ stratified sampling, incorporating equal representation from five street scene categories identified through a $k$-means clustering algorithm. From each category, 100 images were randomly selected, yielding a total of 500 images. We invited 176 participants to perform 18 pairwise comparisons for each of the 19 VPI and VATA indicators. On average, each indicator was compared 6.34 times per image, resulting in 3168 pairwise comparison results per indicator. The methodology is detailed in Section 3.2, with relevant visuals provided in Figures 4 and 5.
A sample size of 500 images successfully captures the diversity of street scene features, as demonstrated in Figure 15. Subfigure (a) highlights that a sample of 500 images captures the overall distribution of street scenes across the first two principal components (PC1 and PC2) with a similarity of 0.95, ensuring comprehensive representation. Beyond this size, the marginal benefit of additional samples diminishes significantly. Stratified sampling further improved coverage of the feature space, increasing it from 0.47 to 0.67. This performance matches what would typically require random sampling of approximately 4,000 images. Subfigure (b) visualizes this improvement, showing that the 500 sampled images (red dots) uniformly represent the overall distribution (blue dots) across the first 11 principal components.
Figure 15 Subfigures (c) and (d) demonstrate that a training sample size of 300 images is sufficient to estimate the average VATA score within a 95% confidence interval (CI), with an error margin of only 5%, while 200 images from the survey sample are allocated for use in the validation and test sets. This estimation assumes an ideal population mean of 2.5 (based on the normal distribution assumption of the TrueSkill algorithm) and a data-augmented population standard deviation of 1. Additionally, increasing the sample size does not result in a significant reduction in CI width, as shown by the graph illustrating the second derivative of the CI width. Moreover, as the sample size increases from 300 to 500, the distribution of VATA scores remains largely unchanged (K-S statistic $<$ 0.05), further suggesting that a representative and convergent sample of 300 images is convergent and sufficient for training a VATA prediction model.
Figure 15: Rationale analytics of survey sample size (500) and training sample size (300)
@@ Appendix E Details on MTNNL model training data and process
Figures 16.a and 16.c demonstrate that different streetscape categories can influence their VATA scores obtained from the online survey. During the division of train/validation/test sets, it is ensured that an equal quantity of each of the five streetscape categories is included in every fold. Figure 16.b illustrates the training process for fold 1 in k-fold cross-validation model training. Both validation loss and train loss consistently decrease and stabilise, indicating that no overfitting occurred during training.
@@ Appendix F Details on stage-1 ENRM models
From the stage-1 ENRM models shown in Figure 6, we can observe that the $R^{2}$ values for IF-VPI inference modelling in the microclimate inference and environmental assessment sections—except for ‘material comfort’—exceed 0.6. This suggests that these VPIs are highly predictable and can be effectively modelled using IFs. Additionally, the VPIs ‘enclosure’, ‘transparency’, and ‘complexity’ in the design quality section demonstrate strong consistency, while in the evoked emotion section, only the $R^{2}$ for ‘lively’ exceeds 0.6. This indicates that most VPIs related to evoked emotions vary significantly between individuals. It is crucial to consider both IF and VPI dimensions in VATA evaluation, as visual-perceptual indicators capture human visual-perceptual interpretations that are not fully reflected in SVI image features.
@@
Appendix G Feature weight analytics in the ENRM inference modelling
Figure 17: Sankey diagram of positive feature weight proportion in the IF-VPI-VATA two-stage inference modelling.
Figure 18: Sankey diagram of negative feature weight proportion in the IF-VPI-VATA two-stage inference modelling
Appendix H Further analytics on feature-wise relationships
H.1 Correlation among VATA and visual-perceptual indicators (VPI)
@@
H.2 Relationship between image features (IF) and visual-perceptual indicators (VPI)